Find dy.



3. sin(xy) = x ^ 2 - y

1
Expert's answer
2021-11-26T00:02:25-0500

"\displaystyle \sin(xy)=x^2 -y\\
\text{Differentiating implicitly, we have}\\
y\cos(xy)+x\cos(xy)\frac{dy}{dx}= 2x-\frac{dy}{dx}\\
\text{Collecting like terms, we have }\\
\frac{dy}{dx}(x\cos(xy)+1)= 2x-y\cos(xy)\\
\implies \frac{dy}{dx} = \frac{2x-y\cos(xy)}{x\cos(xy)+1}"